# Enabling Physical Localization of Uncooperative Cellular Devices

### v2

@@ 4.2.2 Design of Scheduling Manipulation Attack
Our approach exploits BSR [48] and scheduling requests to induce unintended uplink transmissions and maintain an unexpired radio connection. Scheduling requests and BSR, governed by the PHY and MAC layers, lack security protection such as encryption or integrity checks. The core idea is to manipulate uplink scheduling by falsifying the victim’s uplink channel, prompting the eNB to allocate uplink grants to the victim, even when there is no actual uplink data to transmit. Our approach unfolds through three distinct steps, as illustrated in Figure 5.
@@
1) Acquiring user-dedicated radio configuration. The UMA attacker aims to identify the victim’s user-dedicated radio configuration. The user-dedicated radio configuration, which includes the schedulingRequestConfig, is communicated to each UE by the eNB during the establishment of a radio connection. This information is conveyed through the RRC Connection Setup message. The attacker can discern the victim UE’s configuration by monitoring the RRC Connection Setup when the victim attempts to establish a radio connection with the eNB.
@@
2) Establishing forged uplink channel. The UMA attacker manipulates the victim’s uplink channel by utilizing the acquired radio configuration and RNTI of the victim. Initially, the attacker establishes a radio connection with the eNodeB to create a communication channel. Subsequently, the attacker configures its RNTI, and schedulingRequestConfig [2, 5] with identical values to those of the victim. This setting allows the attacker to gain control of the victim’s communication channel.
3) Manipulating uplink scheduling. After acquiring the victim’s communication channel, UMA attacker injects fake uplink scheduling request and fake BSR to the eNB to manipulate uplink scheduling. First, a fake scheduling request is transmitted over the victim’s resource on PUCCH, which is obtained by forging the victim’s uplink channel. Then, the eNB allocates uplink resources to the victim’s RNTI on PUSCH. Second, to obtain multiple uplink grants, the attacker transmits fake BSR over the allocated resource through fake scheduling request. The fake BSR consists of a) the victim’s RNTI, b) the required buffer size, and c) the Logical Channel ID (LCID). In this attack, we set the required buffer size to 200 bytes. The LCID has a value of 3 which indicates data bearer communication in uplink logical channels. The attacker repeatably conducts the above process to make the eNB continuously allocate uplink grants to the victim. At this time, manipulated scheduling requests and BSR are transmitted only using 10% of allocated uplink resources to minimize the impact on the victim’s Quality of Service (QoS).
4.3 Boosting Victim’s Uplink Signal
@@
In this section, we present a novel approach called Power Boosting Attack to enforce a UE to increase its uplink transmission power. This attack is required to overcome a shadow area where the victim’s signals are undetectable to the localization attacker and the impact of a cellular repeater. We first introduce a mechanism for the eNB to control the UE’s uplink transmission power (§ 4.3.1). We then present the implementation of the power boosting attack exploiting TPC commands (§ 4.3.2).
4.3.1 Transmit Power Control
TABLE II: $\Delta_{PUSCH}$ according to TPC command in DCI 0.
@@ 4.3.1 Transmit Power Control
An eNB is responsible for all aspects of data transmission with the connected UEs. Along with the data transmission scheduling, it also controls the data transmission power of the UE by Transmit Power Control (TPC) command, contained in DCI 0 messages. The UE adjusts its transmission power based on the internal power control algorithm, which uses a designated value in the TPC command along with the current self-monitored channel quality [2]. One widely used algorithm adopts the accumulated power control that gradually adjusts the transmission power according to TPC commands as defined in Table II, with a maximum power of 23dBm.
4.3.2 Design of the Power Boosting Attack
Our novel approach to changing the target UE’s uplink transmission power without any privilege is to utilize this TPC command. The TPC commands are delivered over DCI 0 messages without any security protection (i.e., encryption or integrity check). In other words, it is exposed to a downlink overshadowing attack [54]. The key idea is to inject manipulated subframes that contain DCI 0 broadcast to the victim UE, whose TPC is set to increase the victim UE’s transmission power. This power boosting attack can be realized by following three concrete steps.
1) Acquiring configuration for radio communication. To identify the structure of a legitimate downlink subframe, the UMA attacker retrieves the physical configuration of the target eNB. This configuration includes target Physical Cell Identification (PCI), channel bandwidth, Physical channel HybridARQ Indicator Channel (PHICH) configuration, cyclic prefix, and transmission mode of the eNB. This can be retrieved by decoding the messages that the eNB broadcasts such as Master Information Block (MIB), Primary Synchronization Signal (PSS), and Secondary Synchronization Signal (SSS).
@@ 4.3.2 Design of the Power Boosting Attack
3) Overshadowing the manipulated subframe. By extending the signal overshadowing attack [54], UMA attacker injects the crafted subframe with the proper subframe number over PDCCH. More precisely, we mainly modify the subframe construction logic of the SigOver attack and use almost the same logic for injecting the resulting subframe. In this work, we inject the manipulated subframe to overshadow every legitimate subframe that has index number 9. This is mainly because 1) we want to minimize the impact on the victim’s QoS, and 2) there is a race condition with the eNB. As the eNB also tries to adjust the UE’s uplink transmission power, the UE decreases its power whenever it receives the DCI 0 messages from the eNB through the remaining subframes. Therefore, we repeatedly conduct the overshadowing attack with the manipulated subframe.
4.3.3 Strategy for defeating cellular repeater
UMA could distinguish the victim’s signal from the repeater’s signal by employing a power boosting attack. The key idea is to exploit the cellular repeater’s operational logic, specifically designed to amplify the input signal with its maximum output power [^1] The maximum transmission power of the repeater is typically specified in its product documentation., enhancing the overall LTE channel quality. Thus, even after subjecting the victim UE to a power boosting attack, the uplink signal amplified by the repeater remains unchanged.
In this context, the attacker conducts the following procedures. First, the attacker injects a manipulated TPC command having a value of 3 over DCI 0 message. Second, both the victim UE and the repeater receive that message. Third, the victim UE increases its uplink signal strength, but the repeater does not, as it already transmits the uplink signal with its highest power. As a result, it leads to an increase in the arrival signal strength only on the victim UE side, as observed from the sniffer’s perspective. Based on this strategy, the UMA attacker can correctly determine the victim UE’s direction, overcoming the practical challenge posed by the presence of the cellular repeater.
This strategy is viable due to the distinct operational logic of the User Equipment (UE) and the repeater. The repeater serves as a passive network component, solely amplifying the input signal with its maximum output power before relaying the amplified signal. In contrast, the UE dynamically adjusts its uplink transmission power based on the signal path loss, a topic discussed in § 3.3. Consequently, when subjected to a power boosting attack, the UE and repeater exhibit different responses. The repeater consistently amplifies the signal with its maximum power output, while the UE adapts its transmission power dynamically. This discrepancy in the reactions to power boosting allows the attacker to distinguish and identify the victim’s signal effectively.
TABLE III: Test environments for each experiment.
@@ 4.3.3 Strategy for defeating cellular repeater
RNTI acquisition
ON (§ 5.1)
Scheduling manipulation attack
LT (§ 5.2), CT (§ 5.4)
Power boosting attack
LT (§ 5.3), CT (§ 5.4)
End-to-end evaluation of UMA
LT (§ 5.5)
Defeating cellular repeater
ON (§ 5.6)
Iterative direction finding of a UE
ON (§ 5.7)
Multiangulation using uplink signals
ON (§ B.2)
Existence of shadow area
ON (§ B.3)
Received strengths for LTE services
ON (§ B.4)
  • ON: Operational network, LT: Lab testbed using srsRAN, CT: Commercial testbed
TABLE IV: COTS UE used in the experiments.
@@
iPhone $*{}^{*}$ Intel PMB 9955 S -
Redmi Note 9T $*{}^{*}$ Mediatek Dimensity 800U S -
  • S: Scheduling manipulation attack (§ 5.2 and § 5.4), P: Power boosting attack (§ 5.3 and § 5.4)
@@
  • *** Power boosting attack could not be confirmed on these devices due to the incompatibility of their baseband chipsets with our XCAL [53], which is required to monitor uplink transmission power. However, given that their baseband chipsets are also designed to follow 3GPP standards, it is expected that power boosting attack should work on these devices.
5 Evaluation of UMA
@@
We demonstrate the UMA in three different testing environments; a lab testbed (LT), a commercial testbed (CT), and an operating commercial network (ON). Table III summarizes conducted experiments. First, we validate the RNTI acquisition (§ 5.1) in a busy commercial network, affirming the feasibility of scheduling manipulation (§ 5.2) and power boosting (§ 5.3) in the lab testbed. Subsequently, we demonstrate scheduling manipulation and power boosting attacks in the commercial testbed, highlighting their practicality (§ 5.4). An end-to-end evaluation of UMA is then presented, showcasing how the proposed approaches can be integrated to perform the physical localization of uncooperative cellular devices (§ 5.5). Subsequently, we show that power boosting attack can be used to distinguish the UE from the cellular repeater in the commercial network (§ 5.6). Finally, we provide a real-world scenario demonstration where the adversary tracks down the target from outside the building to the front door in an urban commercial network through iterative direction finding (§ 5.7).
@@
Experimental environments. The lab testbed is built using an open-source LTE platform, srsRAN [47], supporting a full-stack LTE eNB. The commercial testbed, established by a national institute, comprises an industry-grade LTE solution from Nokia, situated in a shielded room measuring 3m $\times$ 10m. In addition, we used the operational commercial network of a mobile carrier with the largest subscriber base in [our country]. We evaluated scheduling manipulation and power boosting using nine COTS devices (Table IV) in testbed environments (Table III), due to ethical considerations. Both attacks were implemented by modifying srsRAN [47] with USRP X310 [51], following the aforementioned steps in § 4.2.2 and § 4.3.2.
5.1 RNTI Acquisition
@@
To validate the proposed RNTI acquisition procedure, we assess its ability to identify the victim’s RNTI based on their online identity, even in a crowded operational network.
@@
Initially, we investigate the existence of a distinguishable voice/SMS traffic pattern. To do so, we build a dataset comprising over 4 hours of downlink traffic from two operational LTE networks in two countries. This dataset involves 10–60 active users connected to the eNB and a total of 2,142 radio connections. We could observe that, as an example case, none of the users has a traffic pattern featuring multiple messages/calls with a time gap of over 6 seconds. This observation suggests that the attacker could identify the target victim’s RNTI by carefully designing a transmission strategy for the distinctive traffic pattern.
@@
With the identified traffic pattern and the target’s online identity, we execute the RNTI acquisition process in a crowded commercial network with 862 active users detected over an hour. Using an iPhone 14 Pro as the sender and a Samsung Galaxy S10 5G as the receiver, we conduct this trial ten times. We could confirm that the connection identified by the target’s RNTI is distinguishable from others, achieving a 100% success rate over all trials.
5.2 Scheduling Manipulation Attack
@@
We investigate the effectiveness of manipulating uplink scheduling in UMA. To demonstrate the feasibility of the scheduling manipulation attack, we conduct evaluations using nine COTS devices in the testbed environment (Table IV). To prevent the UE from requesting uplink grants during the attack, we disable all applications, including background processing.
@@
First, we assess whether the victim’s UE generates uplink traffic even when no data is available for transmission, and uplink grants are allocated by the eNB. To inspect the UE’s uplink transmission, we collect uplink packets of the MAC layer at srsENB during scheduling manipulation. Next, we investigate whether the victim’s RNTI remains unchanged during the manipulation of uplink scheduling. Throughout the attack, we monitor the radio status of the UE on both the srsENB and UE sides.
@@
Surprisingly, we verify that the UE transmits uplink traffic to the eNB whenever uplink grants are allocated, even when these grants are assigned through manipulated scheduling requests and BSR (and it has no data to transmit in its buffer). Additionally, through the scheduling manipulation attack, the radio connection between the UE and eNB is consistently maintained and the RNTI of the target UE remains unchanged.
5.3 Power Boosting Attack
We investigate whether the TPC injection indeed increases UE’s transmission power up to the maximum level. To demonstrate the effectiveness of the power boosting attack in UMA, we run the experiment in the testbed environment, using six COTS devices, as described in Table IV.

### v3

@@ 4.2.2. Design of Scheduling Manipulation Attack
Our approach exploits BSR (Tan et al., 2021) and scheduling requests, which manipulate the victim’s uplink scheduling procedure. Our work makes use of them to induce unintended uplink transmissions and maintain an unexpired radio connection. Scheduling requests and BSR, governed by the PHY and MAC layers, lack security protection such as encryption or integrity checks. The core idea is to manipulate uplink scheduling by falsifying the victim’s uplink channel, prompting the eNB to allocate uplink grants to the victim, even when the victim has no actual uplink data to transmit. Our approach unfolds through four distinct steps, as illustrated in Figure 4.
@@
1) Acquiring victim’s radio configuration. The UMA attacker seeks to identify the victim’s user-dedicated radio configuration, including schedulingRequestConfig. This configuration is sent to each UE by the eNB during radio connection establishment via the RRC Connection Setup message. As this signaling is unencrypted, the attacker can discern the victim UE’s configuration by monitoring the downlink channel when the victim initiates a radio connection with the eNB. From this intercepted data, the attacker can extract the victim’s schedulingRequestConfig associated with its RNTI.
@@
2) Establishing forged uplink channel. The UMA attacker forges the victim’s uplink channel using the acquired radio configuration and RNTI of the victim. Initially, the attacker establishes a radio connection with the eNB to create a communication channel. Subsequently, the attacker configures its RNTI and schedulingRequestConfig with identical values to those of the victim. This setting allows the attacker to gain control of the victim’s communication channel. Consequently, the attacker impersonates the victim UE, while the victim maintains its legitimate connection.
3) Manipulating scheduling request. After acquiring the victim’s communication channel, the UMA attacker injects a fake scheduling request to the eNB over the victim’s resource on PUCCH, which is obtained by forging the victim’s uplink channel. Since the scheduling request, indicating the UE has data to transmit, is not authenticated, the attacker can forge this request. Subsequently, the eNB allocates uplink resources to the victim’s RNTI on PUSCH, informing the victim UE it can send packets and report its buffer status.
4) Reporting fake buffer status. To obtain multiple grants, the attacker transmits a packet to the eNB over the allocated resource through the fake scheduling request, containing arbitrary data and a fabricated BSR. The fabricated BSR consists of the required buffer size for future transmission and the Logical Channel ID (LCID). In this attack, we set the required buffer size to 200 bytes. The LCID has a value of 3 which indicates data bearer communication in uplink logical channels. Simultaneously, the victim UE transmits a packet with a BSR informing no pending data. Both transmissions occur in the same subframe, but the attacker’s signal overshadows the victim’s due to the capture effect (Yang et al., 2019; Erni et al., 2022). The eNB then issues multiple uplink grants to the victim’s RNTI based on the fabricated BSR. In response to these unexpected grants, the victim UE transmits dummy packets containing padding data and a BSR indicating no further data to transmit.
The attacker repeatedly conducts the above process to have the eNB continuously allocate uplink grants to the victim. During this process, manipulated messages are transmitted only using 10% of allocated uplink resources to minimize the impact on the victim’s Quality of Service (QoS).
4.3. Boosting Victim’s Uplink Signal
@@
We present a novel approach called Power Boosting Attack to enforce a UE to increase its uplink transmission power. This attack is required to overcome a shadow area where the victim’s signals are undetectable to the localization attacker and the impact of a cellular repeater. We first introduce a mechanism for the eNB to control the UE’s uplink transmission power (§ 4.3.1). We then present the implementation of the power boosting attack exploiting the mechanism (§ 4.3.2).
4.3.1. Transmit Power Control
Table 2. $\Delta_{PUSCH}$ according to TPC command in DCI 0.
@@ 4.3.1. Transmit Power Control
An eNB is responsible for all aspects of data transmission with the connected UEs. Along with the data transmission scheduling, it also controls the data transmission power of the UE by Transmit Power Control (TPC) command, contained in DCI 0 messages. The UE adjusts its transmission power based on the internal power control algorithm, which uses a designated value in the TPC command along with the current self-monitored channel quality (3GPP. TS 36.213, v17.1.0, 2022). One widely used algorithm adopts the accumulated power control that gradually adjusts the transmission power according to TPC commands as defined in Table 2.
4.3.2. Design of Power Boosting Attack
Our novel approach to changing the target UE’s uplink transmission power without any privilege utilizes TPC command. TPC commands are delivered via DCI 0 messages without any security protection (neither encryption nor integrity checks). This leaves them exposed to downlink overshadowing attack (Yang et al., 2019). The key idea is to inject manipulated subframes containing DCI 0 broadcasts to the victim UE, with TPC set to increase the victim UE’s transmission power. This power boosting attack can be executed by following steps.
1) Acquiring configuration for radio communication. To identify the structure of a legitimate downlink subframe, the UMA attacker retrieves the physical configuration of the target eNB. This configuration includes target Physical Cell Identification (PCI), channel bandwidth, Physical channel HybridARQ Indicator Channel (PHICH) configuration, cyclic prefix, and transmission mode of the eNB. This can be retrieved by decoding messages that the eNB broadcasts such as Master Information Block (MIB), Primary Synchronization Signal (PSS), and Secondary Synchronization Signal (SSS).
@@ 4.3.2. Design of Power Boosting Attack
3) Overshadowing manipulated subframe. The UMA attacker extends the signal overshadowing (SigOver) attack (Yang et al., 2019), injecting a crafted subframe with a proper subframe number over the downlink control channel. More precisely, we mainly modify the subframe construction logic of SigOver (Yang et al., 2019), while using nearly the same approach for subframe injection. In this work, we inject a manipulated subframe to overshadow every legitimate subframe with index number 9. This choice serves two purposes: 1) to minimize the impact on the victim’s QoS, and 2) to address a race condition with the eNB. Since the eNB also attempts to adjust the UE’s uplink transmission power, the UE decreases its power upon receiving DCI 0 messages from the eNB. Consequently, we repeatedly conduct the overshadowing attack with the manipulated subframes.
4.3.3. Strategy for defeating cellular repeater
UMA could distinguish the victim’s signal from the repeater’s signal by employing a power boosting attack. The key idea is to exploit the cellular repeater’s operational logic, specifically designed to amplify the input signal with its maximum output power, enhancing the overall LTE channel quality. Thus, even after subjecting the victim UE to the power boosting attack, uplink signals amplified by the repeater remain unchanged.
In this context, the attacker conducts the following procedures. First, the attacker injects a manipulated TPC command having a value of 3 over DCI 0 message. Second, both the victim UE and the repeater receive that message. Third, the victim UE increases its uplink signal strength, but the repeater does not, as it already transmits (relays) the uplink signal with its highest power. As a result, it leads to an increase in the arrival signal strength only on the victim UE side, as observed from the sniffer’s perspective. Based on this strategy, the UMA attacker can correctly determine the victim UE’s direction, overcoming the practical challenge posed by the presence of the cellular repeater.
This strategy is viable due to the distinct operational logic of the UE and repeater. The repeater serves as a passive network component, solely amplifying the input signal with its maximum output power before relaying. In contrast, the UE dynamically adjusts its uplink transmission power based on the signal path loss, a topic discussed in § 3.3. Consequently, when subjected to the power boosting attack, the UE and repeater exhibit different responses. The repeater consistently amplifies the signal with its maximum power output, while the UE adapts its transmission power dynamically. This discrepancy in reactions to power boosting allows the attacker to distinguish and identify the victim’s signal effectively.
Table 3. Test environments for each experiment.
@@ 4.3.3. Strategy for defeating cellular repeater
RNTI acquisition ON (§ 5.1)
Scheduling manipulation attack LT (§ 5.2), CT (§ 5.4)
Power boosting attack LT (§ 5.3), CT (§ 5.4)
End-to-end evaluation of UMA LT (§ 5.5)
Defeating cellular repeater CT (§ 5.6)
Multiangulation using uplink signals ON ((UMA, 2024))
Existence of shadow area ON ((UMA, 2024))
Received strengths for LTE services ON ((UMA, 2024))
  • ON: Operational network, LT: Lab testbed using srsRAN, CT: Commercial testbed.
Table 4. COTS UE used in the experiments.
@@
iPhone XS Intel PMB 9955 S -
Redmi Note 9T 5G Mediatek Dimensity 800U S -
  • S: Scheduling manipulation attack (§ 5.2 and § 5.4), P: Power boosting attack (§ 5.3 and § 5.4).
  • – Only three types of devices were tested on the commercial testbed, due to compatibility issues with the environment.
@@
  • $\ast$ Power boosting attack could not be confirmed on these devices due to the incompatibility of their baseband chipsets with our XCAL (Accuver, 2015), which is required to monitor uplink transmission power. However, given that their baseband chipsets are also designed to follow 3GPP standards, it is expected that power boosting attack should work on these devices.
5. Evaluation of UMA
@@
We demonstrate the UMA in three different testing environments (Table 3): a lab testbed (LT), a commercial testbed (CT), and an operating commercial network (ON). First, we validate the RNTI acquisition (§ 5.1) in a busy commercial network, and then prove scheduling manipulation (§ 5.2) and power boosting (§ 5.3) in the lab testbed. Subsequently, we demonstrate scheduling manipulation and power boosting attacks in the commercial testbed, highlighting their feasibility (§ 5.4). An end-to-end evaluation of UMA is then presented, showcasing how the proposed approaches can be integrated to perform the physical localization of uncooperative cellular devices (§ 5.5). Finally, we show the effectiveness of the power boosting attack to distinguish the UE from the cellular repeater in the commercial testbed (§ 5.6).
@@
Experimental environments. The lab testbed is built using an open-source LTE platform, srsRAN (Systems, 2019). The commercial testbed, established by a national institute, comprises an industry-grade LTE solution from Nokia, situated in a shielded room measuring 5m $\times$ 7m. In addition, we used the operational commercial network of a mobile carrier with the largest subscriber base in South Korea. We evaluated scheduling manipulation and power boosting using nine COTS devices in testbed environments. Both attacks were implemented by modifying srsRAN (Systems, 2019) with USRP X310 (Research, 2014).
5.1. RNTI Acquisition
@@
We first assess its ability to identify the victim’s RNTI based on their online identity, in a crowded operational network.
@@
Initially, we investigate the existence of a distinguishable voice/SMS traffic pattern. To do so, we build a dataset comprising over 4 hours of downlink traffic from two operational LTE networks in two countries. This dataset involves 10–60 active users connected to the eNB and a total of 2,142 radio connections. We could observe that, as an example case, none of the users has a traffic pattern featuring multiple messages/calls with a time gap of over 6 seconds. It suggests that the attacker could identify the target UE’s RNTI by carefully designing a transmission strategy for the distinctive pattern.
@@
With the identified traffic pattern and the target’s online identity, we execute the RNTI acquisition process in a crowded commercial network with 862 active users detected over an hour. Using an iPhone 14 Pro (sender) and a Galaxy S10 5G (receiver), we conduct ten trials. We could confirm that the connection identified by the target’s RNTI is distinguishable, achieving a 100% success rate over all trials.
5.2. Scheduling Manipulation Attack
@@
We evaluate the feasibility of the scheduling manipulation attack. For this, we use nine COTS devices in the testbed environment (Table 4). To prevent the UE from requesting uplink grants during the attack, we disable all applications.
@@
First, we assess whether a) the victim’s UE generates uplink traffic even when no data is available for transmission and b) uplink grants are allocated by the eNB. To inspect the UE’s uplink transmission, we collect uplink packets of the MAC layer at eNB during scheduling manipulation. Next, we investigate whether the victim’s RNTI remains unchanged during the manipulation of uplink scheduling and monitor the radio status of the UE on both the srsENB and UE sides.
@@
Surprisingly, we verify that the UE transmits uplink traffic to the eNB whenever uplink grants are allocated, even when these grants are assigned through manipulated scheduling requests and BSR (and it has no data to transmit in its buffer). Additionally, through the scheduling manipulation attack, the radio connection between the UE and eNB is consistently maintained and the target UE’s RNTI remains unchanged.
Our website (UMA, 2024) provides packet capture images during scheduling manipulation. These captures illustrate the result of executing the attack four times consecutively, including the contents of both the attacker’s and the target’s packets.
We further investigated the root causes of continuous uplink transmission even when the UE has no data to transmit in its current buffer. We examine the memory status by dumping the Communication Processor (CP) called RAMDUMP (Golde and Komaromy, 2016), and confirm that uninitialized data (i.e., previously transmitted buffer) in CP memory is transmitted.
5.3. Power Boosting Attack